a magnetic field of 3.6T is applied to a paramagnetic gas. The atoms of the gas have magnetic dipole moment of 4.5*10^-23 J/T. At what temperature will the mean transitional kinetic energy of an atom of the gas equal to the energy required to change the alignment of atoms magnetic dipole from antiparallel to parallel(to the magnetic field)

Grade:12

1 Answers

Arun
5 years ago
Dear student,Energy required to change orientation=NMB (- cos 180 – cos 0) = 2NMB = 10.8 * 10^-23 This is equal to mean translational kinetic energy=3/2 Nk T = 10.8 * 10^-23 
 
Now find T